# Cascaded Large-Scale TSP Solving with Unified Neural Guidance: Bridging Local and Population-based Search

### v1

@@ Abstract
The traveling salesman problem (TSP) is a fundamental NP-hard optimization problem. This work presents UNiCS, a novel unified neural-guided cascaded solver for solving large-scale TSP instances. UNiCS comprises a local search (LS) phase and a population-based search (PBS) phase, both guided by a learning component called unified neural guidance (UNG). Specifically, UNG guides solution generation across both phases and determines appropriate phase transition timing to effectively combine the complementary strengths of LS and PBS. While trained only on simple distributions with relatively small-scale TSP instances, UNiCS generalizes effectively to challenging TSP benchmarks containing much larger instances (10,000-71,009 nodes) with diverse node distributions entirely unseen during training. Experimental results on the large-scale TSP instances demonstrate that UNiCS consistently outperforms state-of-the-art methods, with its advantage remaining consistent across various runtime budgets.
@@ 1 Introduction
The traveling salesman problem (TSP) is a fundamental NP-hard optimization problem with extensive applications Hubert and Baker (1978); Bland and Shallcross (1989); Aoyama et al. (2004). Given a set of nodes and the distances between them, the goal of TSP is to find the shortest possible tour that visits each node exactly once and returns to the starting node. Over the past decades, plenty of methods have been proposed to solve TSP. Exact methods, including highly optimized ones like the Concorde solver Applegate et al. (2006), are guaranteed to find optimal solutions but suffer from worst-case exponential time complexity. This limitation makes them impractical for solving large-scale TSP instances. In contrast, heuristic search methods Helsgaun (2017), although without guaranteeing optimality, can find high-quality solutions within reasonable computational time.
@@
In general, most (if not all) traditional heuristic methods are designed based on expert knowledge, making them human-interpretable. However, with the advancement of deep learning technologies over the past decade, there has been growing interest in training powerful deep neural networks to solve TSP instances generated from specific distributions Vinyals et al. (2015); Kool et al. (2019). Unfortunately, despite continuous improvements in performance, these deep learning models still fall significantly behind traditional heuristic methods in terms of solution quality Liu et al. (2023). This performance gap becomes particularly evident when dealing with large-scale TSP instances containing thousands of nodes and real-world TSP instances that may not conform to the distributions used during training.
@@
Rather than training purely neural networks to solve TSP, combining strong traditional heuristic methods with learning techniques is believed to be a more practical approach for advancing the state-of-the-art TSP solving capabilities, as highlighted by Bengio et al. (2021). Through learning from experience, the learning components, which are integrated into the heuristic methods, aim to discover better policies to replace hand-crafted rules for making critical algorithmic decisions during the search process. Throughout this paper, these methods are refereed to as hybrid methods. Representative hybrid methods for solving TSP include NLKH Xin et al. (2021) and VSR-LKH Zheng et al. (2023a) that enhance the powerful Lin-Kernighan-Helsgaun (LKH) method Helsgaun (2017) by learning policies for candidate edge generation and selection during LKH’s $\lambda$-opt search process, respectively. Both NLKH and VSR-LKH have shown considerable improvements over the original LKH method across a wide range of TSP instances. However, they are still inherently limited by the underlying local search (LS) mechanism of LKH. In general, LS methods often exhibit limited global search capabilities and tend to get trapped in local optima. This becomes particularly problematic for large-scale TSP instances, where the presence of numerous local optima can significantly impact the solution quality obtained by LS methods such as LKH and its hybrid variants.
@@
Besides LS, population-based search (PBS) like genetic algorithm (GA) represents another powerful heuristic framework for solving TSP. By maintaining and updating a population of solutions, PBS typically exhibits stronger capabilities in exploring the solution space compared to LS methods. Taking advantage of this, powerful PBS methods such as GA with edge assembly crossover (EAX) Nagata and Kobayashi (2013a) and multi-agent optimization system (MAOS) Xie and Liu (2009) can achieve better solution quality than LS methods like LKH, when solving TSP instances with more than 10,000 nodes. However, due to the computational overhead from evolving many solutions simultaneously, PBS methods suffer from slower convergence – they cannot yield high-quality solutions as rapidly as LS methods.
@@
Unlike many existing deep learning models that focus on solving TSP instances from specific distributions, UNiCS is evaluated under more challenging settings. Trained only on simple distributions with relatively small-scale instances, UNiCS is tested on conventional TSP benchmarks TSPlib Reinelt (1991), National Cook (2013), and VLSI Rohe (2013) that differ substantially from the training data. These benchmarks contain much larger instances (with 10,000 to 71,009 nodes) and feature diverse node distributions entirely unseen during training. Experimental results on these instances show that UNiCS achieves superior solution quality compared to state-of-the-art heuristic methods and hybrid methods with learning components. Moreover, the performance advantage of UNiCS remains consistent under various runtime budgets, demonstrating its strong practicality to different real-world time constraints.
@@ 2.1 Traditional Methods
Traditional methods for solving TSP can be broadly classified into exact and heuristic ones. Exact methods like Concorde Applegate et al. (2007) utilize branch-and-cut techniques to find optimal solutions, but are impractical for large-scale instances due to exponential time complexity. Heuristic methods, on the other hand, aim to find near-optimal solutions efficiently. These can be further divided into LS-based and PBS-based ones Gutin and Punnen (2006); Xie and Liu (2009), with LKH Helsgaun (2017) and EAX Nagata and Kobayashi (2013a) being representative methods of each category respectively. LKH adopts iterative local search with $\lambda$-opt moves and $\alpha$-nearness measure to reduce search space, offering quick high-quality solutions but tends to get trapped in local optima. EAX maintains a population of solutions and evolves them through sophisticated edge assembly crossover, showing strong performance particularly on large-scale instances. Since EAX forms the foundation for the PBS phase in UNiCS, its key components are introduced below (see Appendix C for a detailed description of EAX). While LKH similarly forms the foundation for the LS phase in UNiCS, its implementation under UNG aligns with NLKH and can be found in Appendix D.
@@
[…]
@@
The $AB$-cycle selected in step 2 directly determines the $E$-set and the resulting offspring, making it the crucial part of the EAX method. Indeed, various selection strategies were designed and evaluated in Nagata and Kobayashi (2013a), and the final adopted approach in EAX was random selection in Stage I and Tabu search in Stage II. In the PBS phase of our proposed UNiCS, a learning component (UNG) replaces these hand-crafted rules to make better $AB$-cycle selections.
@@ 2.2 Deep Learning Models
Deep learning models for solving TSP can be categorized into constructive and improvement ones. The former build solutions sequentially, starting with Pointer Network Vinyals et al. (2015) and advancing through reinforcement learning Bello et al. (2017) and Transformer-based architectures Kool et al. (2019). Recent works have explored employing solution symmetry Kwon et al. (2020), policy ensemble Gao et al. (2023), and divide-and-conquer strategies Pan et al. (2023); Ye et al. (2024) to tackle larger TSP instances. Improvement models focus on improving existing solutions, starting with controlling 2-opt operator and advancing to novel node embedding Ma et al. (2021) and node pair selection Wu et al. (2022). Despite the ongoing progress, deep learning models still significantly lag behind heuristic methods in solution quality, especially for large-scale TSP instances. For more details on this direction, one may refer to the recent survey Cappart et al. (2023).
@@ 2.3 Hybrid Methods
Hybrid solvers combining heuristic methods with learning techniques have shown promise for large-scale TSP instances. VSR-LKH Zheng et al. (2023a) used reinforcement learning for edge selection during the $\lambda$-opt process of LKH, while NLKH Xin et al. (2021) employed a graph neural network to quickly generate high-quality candidate edge sets for LKH. Built upon VSR-LKH, RHGA Zheng et al. (2023b) combined it with EAX, using the learned Q-values to guide tour generation in EAX. Beyond LKH-based methods, enhancing other types of heuristic methods such as the ant colony optimizer Ye et al. (2023) has also been explored. Another research direction focuses on incorporating different TSP heuristic methods into algorithm portfolios, where methods either run independently in parallel Liu et al. (2019); Tang et al. (2021); Liu et al. (2022) or are selected based on machine learning models Kerschke et al. (2018); Zhao et al. (2021) when solving problem instances. Our proposed solver, UNiCS, differs from existing hybrid solvers in that UNG directs and coordinates both LS and PBS phases throughout the solving process, which is the first attempt of such integration in the literature.
@@ 3.1 Solution Generation Guided by UNG
UNG contains a neural network that scores edges to estimate their likelihood of appearing in optimal solutions. This scoring network guides solution generation in both LS and PBS phases. For the LS phase, the scoring network guides candidate edge generation during LKH’s $\lambda$-opt process. This implementation aligns with NLKH and is thus omitted here due to space limitation (details provided in Appendix D). For the PBS phase, the scoring network guides $AB$-cycle selection during EAX’s crossover process, which is detailed below.
Scoring Network. Inspired by NLKH Xin et al. (2021), the scoring network is implemented as a sparse graph network (SGN) to handle large-scale TSP instances efficiently. The input TSP instance is first converted to a sparse directed graph $G^{*}=(V,E^{*})$, where $V$ represents the set of nodes and $E^{*}$ contains only the $\gamma$-shortest edges originating from each node. The SGN consists of an encoder that embeds edges and nodes into feature vectors, as well as a decoder used for edge scoring. The encoder first linearly projects the node input $x_{v}\in V$ and the edge input $x_{e}\in E^{*}$ into feature vectors $v^{0}_{i}\in R^{D}$ and $e^{0}_{i,j}\in R^{D}$, where $D$ is the feature dimension; then the node and edge features are recursively embedded with $L$-layer SGN:
$\small a_{i,j}^{l}=\exp\left(W_{a}^{l}e_{i,j}^{l-1}\right)\oslash\sum_{(i,k)\in E^{*}}\exp\left(W_{a}^{l}e_{i,k}^{l-1}\right),$ (1)
$\small v_{i}^{l}=\mathcal{F}\left(W_{s}^{l}v_{i}^{l-1}+\sum_{(i,j)\in E^{*}}a_{i,j}^{l}\odot W_{n}^{l}v_{j}^{l-1}\right)+v_{i}^{l-1},$ (2)
$\small r_{i,j}^{l}=\begin{cases}W_{r}^{l}e_{j,i}^{l-1},&\text{if }(j,i)\in E^{*}\\ W_{r}^{l}p,&\text{otherwise}\end{cases},$ (3)
$\small e_{i,j}^{l}=\mathcal{F}\left(W_{f}^{l}v_{i}^{l-1}+W_{t}^{l}v_{j}^{l-1}+W_{o}^{l}e_{i,j}^{l-1}+r_{i,j}^{l}\right)+e_{i,j}^{l-1}.$ (4)
@@
$\small\beta_{i,j}=\frac{\exp\left(W_{\beta}e_{i,j}^{F}\right)}{\sum_{(i,k)\in E^{*}}\exp\left(W_{\beta}e_{i,k}^{F}\right)},$ (5)
@@
$\small\text{score}_{AB}=\sum_{(i,j)\in E_{B}\cap E_{AB}}\beta_{i,j}-\sum_{(i,j)\in E_{A}\cap E_{AB}}\beta_{i,j},$ (7)
@@
3.2 Phase Transition Guided by UNG
On the other hand, it has been widely observed that the complementary strengths of LS and PBS manifest distinctly across TSP instances of different sizes Nagata and Kobayashi (2013a); Zhao et al. (2021). This observation suggests that problem size could serve as a crucial factor in determining the appropriate $t_{trans}$. Based on this insight, a transition policy mapping from problem size to $t_{trans}$ is established within the UNG module. The policy takes the sparse graph $G^{*}$ of a TSP instance as input and outputs the corresponding $t_{trans}$. To train this policy, random uniform TSP instances of various sizes are used as training instances, with UNiCS being run on them under different $t_{trans}$ settings to gather training data (detailed in Section 4.1). For each training instance, different $t_{trans}$ settings are compared by evaluating the area under the optimality gap curve. This metric provides a comprehensive assessment of UNiCS’s performance across the entire runtime rather than at specific runtime points, and its aggregated nature also helps mitigate the impact of randomness in performance assessment. Then, the optimal $t_{trans}$ with the smallest area is identified. After obtaining training data, a linear model is fitted to these optimal values to derive the transition policy. This simple model is chosen to avoid overfitting while maintaining interpretability and computational efficiency. UNG then uses the trained policy to automatically determine $t_{trans}$ for testing instances.
@@ 4 Computational Studies
4.1 Experimental Setup
Training UNG. The scoring network SGN within UNG was trained on two-dimensional Euclidean TSP instances with 500 to 1000 nodes, where node coordinates were uniformly sampled from a unit square. Since the number of optimal edges for each instance is proportional to the number of nodes, $\frac{500000}{V}$ instances were generated for each problem size $V$ in the training set, resulting in approximately 346,000 training instances in total. The optimal edges $E^{*}_{o}$ for each instance were obtained using Concorde. Each instance was converted to a sparse directed graph with $\gamma=20$ edges per node (see Section 3.1). The SGN consists of $L=25$ sparse graph convolution layers with a hidden dimension $D=128$ and 4 attention heads. The model was trained for 100 epochs using Adam optimizer with a learning rate of $10^{-4}$. For the transition policy within UNG, 56 TSP instances ranging from 3,000 to 30,000 nodes were generated using the same uniform distribution described above. For each instance, the transition time was varied from 50 to 650 seconds in 50-second increments. The optimal $t_{trans}$ was then determined by evaluating the area under the optimality gap curve (see Appendix B).
Competitors and Hyperparameter Settings. UNiCS was compared against state-of-the-art heuristic methods EAX Nagata and Kobayashi (2013b) and LKH version 3 Helsgaun (2017), as well as hybrid solvers NLKH Xin et al. (2021), VSR-LKH Zheng et al. (2023a), and RHGA Zheng et al. (2023b). To verify the effectiveness of UNG in guiding $AB$-cycle selection, a version of UNiCS (named UNiCS-P) that removes LS phase and uses only PBS phase was also included in the comparison. Essentially, UNiCS-P is a variant of EAX equipped with UNG guidance. For all EAX-based methods (including EAX and UNiCS-P), the EAX-specific hyperparameters were set according to the original EAX paper. All LKH-based methods (including LKH, NLKH, and VSR-LKH) utilized LKH’s default hyperparameter settings. NLKH and UNiCS shared the same SGN architecture and training procedure. For RHGA, hyperparameters were set as specified in the original paper. The probability of random selection of $AB$-cycles in UNiCS, i.e., $\eta$, was set to 0.5 based on our preliminary experiments (see Appendix A.4).
TSP Benchmarks. All two-dimensional Euclidean TSP instances with more than 10,000 nodes were collected from three TSP benchmark sets: TSPLib Reinelt (1991), National Cook (2013), and VLSI Rohe (2013), totaling 38 instances. These benchmarks represent diverse real-world scenarios. VLSI instances are derived from integrated circuit applications, National instances are based on city distances in different countries, while TSPLib includes various sources such as logistics, circuit board drilling, and transistor routing. The collected instances were further divided into two sets: Large set (32 instances, 10,000-40,000 nodes) and Ex-Large set (6 instances, 40,000+ nodes). The Large set was further categorized by source: VLSI (21 instances), National (7 instances), and TSPLib (4 instances). The largest TSP instance considered here contains 71,009 nodes.
Testing Methods. To make fair comparison, the testing of all methods was conducted on the same server with two AMD EPYC 7713 CPUs and 512GB RAM running Ubuntu 20.04, and each method was evaluated using a single CPU core with 10 independent runs per instance. To ensure reproducibility, random seeds started from 42 and were incremented by 60 for each subsequent run. Three runtime budgets were considered in the experiments: short (1800s), medium (3600s), and long (7200s). Due to significantly inferior performance compared to other methods, RHGA’s results are omitted in the paper and presented in Appendix A.2.
Figure 2: Convergence curves averaged over 10 runs.
Table 1: Optimality gap (%) on the Large set (10 runs per instance). “-” in Best / Avg columns indicates no solutions produced in 10 runs / solutions produced only in some runs. Bold indicates best performance. “Total” row summarizes average gaps across Large set.

### v2

@@ Abstract
The traveling salesman problem (TSP) is a fundamental NP-hard optimization problem. Over the past decades, traditional heuristic methods have achieved substantial success in solving TSP, yet their performance, particularly for large-scale instances, remains to be further improved. The advancement of deep learning technologies over the past decade has driven a growing number of attempts to solve TSP by leveraging neural guidance. However, these efforts predominantly focus on small-scale TSP instances, with limited improvements in solving performance for large-scale instances, revealing persistent scalability challenges. This work presents UNiCS, a novel unified neural-guided cascaded solver for solving large-scale TSP instances. UNiCS comprises a local search (LS) phase and a population-based search (PBS) phase, both guided by a learning component called unified neural guidance (UNG). Specifically, UNG guides solution generation across both phases and determines appropriate phase transition timing to effectively combine the complementary strengths of LS and PBS. While trained only on simple distributions with relatively small-scale TSP instances, UNiCS generalizes effectively to challenging TSP benchmarks containing much larger instances (10,000-71,009 nodes) with diverse node distributions entirely unseen during training. Experimental results on the large-scale TSP instances demonstrate that UNiCS consistently outperforms state-of-the-art methods, with its advantage remaining consistent across various runtime budgets.
[…]
9829
@@ 1 Introduction
The traveling salesman problem (TSP) is a fundamental NP-hard optimization problem with extensive applications [13, 6, 1]. Given a set of nodes and the distances between them, the goal of TSP is to find the shortest possible tour that visits each node exactly once and returns to the starting node. Over the past decades, plenty of methods have been proposed to solve TSP. Exact methods, including highly optimized ones like the Concorde solver [2], are guaranteed to find optimal solutions but suffer from worst-case exponential time complexity. This limitation makes them impractical for solving large-scale TSP instances. In contrast, heuristic search methods [11], although without guaranteeing optimality, can find high-quality solutions within reasonable computational time.
@@
In general, most (if not all) traditional heuristic methods are designed based on expert knowledge, making them human-interpretable. However, with the advancement of deep learning technologies over the past decade, there has been growing interest in training powerful deep neural networks to solve TSP instances generated from specific distributions [29, 14]. Unfortunately, despite continuous improvements in performance, these deep learning models still fall significantly behind traditional heuristic methods in terms of solution quality [18]. This performance gap becomes particularly evident when dealing with large-scale TSP instances containing thousands of nodes and real-world TSP instances that may not conform to the distributions used during training.
@@
Rather than training purely neural networks to solve TSP, combining strong traditional heuristic methods with learning techniques is believed to be a more practical approach for advancing the state-of-the-art TSP solving capabilities, as highlighted by [5]. Through learning from experience, the learning components, which are integrated into the heuristic methods, aim to discover better policies to replace hand-crafted rules for making critical algorithmic decisions during the search process. Throughout this paper, these methods are refereed to as hybrid methods. Representative hybrid methods for solving TSP include NLKH [32] and VSR-LKH [36] that enhance the powerful Lin-Kernighan-Helsgaun (LKH) method [11] by learning policies for candidate edge generation and selection during LKH’s $\lambda$-opt search process, respectively. Both NLKH and VSR-LKH have shown considerable improvements over the original LKH method across a wide range of TSP instances. However, they are still inherently limited by the underlying local search (LS) mechanism of LKH. In general, LS methods often exhibit limited global search capabilities and tend to get trapped in local optima. This becomes particularly problematic for large-scale TSP instances, where the presence of numerous local optima can significantly impact the solution quality obtained by LS methods such as LKH and its hybrid variants.
@@
Besides LS, population-based search (PBS) like genetic algorithm (GA) represents another powerful heuristic framework for solving TSP. By maintaining and updating a population of solutions, PBS typically exhibits stronger capabilities in exploring the solution space compared to LS methods. Taking advantage of this, powerful PBS methods such as GA with edge assembly crossover (EAX) [23] and multi-agent optimization system (MAOS) [31] can achieve better solution quality than LS methods like LKH, when solving TSP instances with more than 10,000 nodes. However, due to the computational overhead from evolving many solutions simultaneously, PBS methods suffer from slower convergence – they cannot yield high-quality solutions as rapidly as LS methods.
@@
Unlike many existing deep learning models that focus on solving TSP instances from specific distributions, UNiCS is evaluated under more challenging settings. Trained only on simple distributions with relatively small-scale instances, UNiCS is tested on conventional TSP benchmarks TSPlib [26], National [8], and VLSI [27] that differ substantially from the training data. These benchmarks contain much larger instances (with 10,000 to 71,009 nodes) and feature diverse node distributions entirely unseen during training. Experimental results on these instances show that UNiCS achieves superior solution quality compared to state-of-the-art heuristic methods and hybrid methods with learning components. Moreover, the performance advantage of UNiCS remains consistent under various runtime budgets, demonstrating its strong practicality to different real-world time constraints.
@@ 2.1 Traditional Methods
Traditional methods for solving TSP can be broadly classified into exact and heuristic ones. Exact methods like Concorde [3] utilize branch-and-cut techniques to find optimal solutions, but are impractical for large-scale instances due to exponential time complexity. Heuristic methods, on the other hand, aim to find near-optimal solutions efficiently. These can be further divided into LS-based and PBS-based ones [10, 31], with LKH [11] and EAX [23] being representative methods of each category respectively. LKH adopts iterative local search with $\lambda$-opt moves and $\alpha$-nearness measure to reduce search space, offering quick high-quality solutions but tends to get trapped in local optima. EAX maintains a population of solutions and evolves them through sophisticated edge assembly crossover, showing strong performance particularly on large-scale instances. Since EAX forms the foundation for the PBS phase in UNiCS, its key components are introduced below (see Section C in the supplementary material[20] for a detailed description of EAX). While LKH similarly forms the foundation for the LS phase in UNiCS, its implementation under UNG aligns with NLKH and can be found in Section D in the supplementary material[20].
@@
  • Step 3: An intermediate solution is created by modifying $p_{A}$. This modification involves removing edges in $E_{A}\cap E\text{-set}$ from $E_{A}$ and adding edges in $E_{B}\cap E\text{-set}$ to $E_{A}$. The resulting intermediate solution may contain one or more sub-tours and thus be invalid. These sub-tours are then combined through a specific merging process to create a valid solution.
@@
The $AB$-cycle selected in step 2 directly determines the $E$-set and the resulting offspring, making it the crucial part of the EAX method. Indeed, various selection strategies were designed and evaluated in [23], and the final adopted approach in EAX was random selection in Stage I and Tabu search in Stage II. In the PBS phase of our proposed UNiCS, a learning component (UNG) replaces these hand-crafted rules to make better $AB$-cycle selections.
@@ 2.2 Deep Learning Models
Deep learning models for solving TSP can be categorized into constructive and improvement ones. The former build solutions sequentially, starting with Pointer Network [29] and advancing through reinforcement learning [4] and Transformer-based architectures [14]. Recent works have explored employing solution symmetry [15], policy ensemble [9], and divide-and-conquer strategies [25, 34] to tackle larger TSP instances. Improvement models focus on improving existing solutions, starting with controlling 2-opt operator and advancing to novel node embedding [22] and node pair selection [30]. Despite the ongoing progress, deep learning models still significantly lag behind heuristic methods in solution quality, especially for large-scale TSP instances. For more details on this direction, one may refer to the recent survey [7].
@@ 2.3 Hybrid Methods
Hybrid solvers combining heuristic methods with learning techniques have shown promise for large-scale TSP instances. VSR-LKH [36] used reinforcement learning for edge selection during the $\lambda$-opt process of LKH, while NLKH [32] employed a graph neural network to quickly generate high-quality candidate edge sets for LKH. Built upon VSR-LKH, RHGA [37] combined it with EAX, using the learned Q-values to guide tour generation in EAX. In contrast to UNiCS, RHGA adopts a memetic algorithm framework, where VSR-LKH is applied to further refine the offspring population generated in each iteration of EAX. Beyond LKH-based methods, enhancing other types of heuristic methods such as the ant colony optimizer [33] has also been explored. Another research direction focuses on incorporating different TSP heuristic methods into algorithm portfolios, where methods either run independently in parallel [16, 28, 17] or are selected based on machine learning models [35] when solving problem instances. Our proposed solver, UNiCS, differs from existing hybrid solvers in that UNG directs and coordinates both LS and PBS phases throughout the solving process, which is the first attempt of such integration in the literature.
@@ 3.1 Solution Generation Guided by UNG
UNG contains a neural network that scores edges to estimate their likelihood of appearing in optimal solutions. This scoring network guides solution generation in both LS and PBS phases. For the LS phase, the scoring network guides candidate edge generation during LKH’s $\lambda$-opt process. This implementation aligns with NLKH and is thus omitted here due to space limitation (details provided in Section D in the supplementary material[20]). For the PBS phase, the scoring network guides $AB$-cycle selection during EAX’s crossover process, which is detailed below.
Scoring Network. Inspired by NLKH [32], the scoring network is implemented as a sparse graph network (SGN) to handle large-scale TSP instances efficiently. The input TSP instance is first converted to a sparse directed graph $G^{*}=(V,E^{*})$, where $V$ represents the set of nodes and $E^{*}$ contains only the $\gamma$-shortest edges originating from each node. The SGN consists of an encoder that embeds edges and nodes into feature vectors, as well as a decoder used for edge scoring. The encoder first linearly projects the node input $x_{v}\in V$ and the edge input $x_{e}\in E^{*}$ into feature vectors $v^{0}_{i}\in R^{D}$ and $e^{0}_{i,j}\in R^{D}$, where $D$ is the feature dimension; then the node and edge features are recursively embedded with $L$-layer SGN:
$a_{i,j}^{l}=\exp\left(W_{a}^{l}e_{i,j}^{l-1}\right)\oslash\sum_{(i,k)\in E^{*}}\exp\left(W_{a}^{l}e_{i,k}^{l-1}\right),$ (1)
$v_{i}^{l}=\mathcal{F}\left(W_{s}^{l}v_{i}^{l-1}+\sum_{(i,j)\in E^{*}}a_{i,j}^{l}\odot W_{n}^{l}v_{j}^{l-1}\right)+v_{i}^{l-1},$ (2)
$r_{i,j}^{l}=\begin{cases}W_{r}^{l}e_{j,i}^{l-1},&\text{if }(j,i)\in E^{*}\\ W_{r}^{l}p,&\text{otherwise}\end{cases},$ (3)
$e_{i,j}^{l}=\mathcal{F}\left(W_{f}^{l}v_{i}^{l-1}+W_{t}^{l}v_{j}^{l-1}+W_{o}^{l}e_{i,j}^{l-1}+r_{i,j}^{l}\right)+e_{i,j}^{l-1}.$ (4)
@@
$\beta_{i,j}=\frac{\exp\left(W_{\beta}e_{i,j}^{F}\right)}{\sum_{(i,k)\in E^{*}}\exp\left(W_{\beta}e_{i,k}^{F}\right)},$ (5)
@@
$\text{score}_{AB}=\sum_{(i,j)\in E_{B}\cap E_{AB}}\beta_{i,j}-\sum_{(i,j)\in E_{A}\cap E_{AB}}\beta_{i,j},$ (7)
@@ 4 $s\leftarrow\text{PBS}(w_{in},s,\beta_{i,j},t_{max}-t_{trans})$;
Figure 2: Convergence curves averaged over 10 runs.
3.2 Phase Transition Guided by UNG
On the other hand, it has been widely observed that the complementary strengths of LS and PBS manifest distinctly across TSP instances of different sizes [23, 35]. This observation suggests that problem size could serve as a crucial factor in determining the appropriate $t_{trans}$. Based on this insight, a transition policy mapping from problem size to $t_{trans}$ is established within the UNG module. The policy takes the sparse graph $G^{*}$ of a TSP instance as input and outputs the corresponding $t_{trans}$. To train this policy, random uniform TSP instances of various sizes are used as training instances, with UNiCS being run on them under different $t_{trans}$ settings to gather training data (detailed in Section 4.1). For each training instance, different $t_{trans}$ settings are compared by evaluating the area under the optimality gap curve (detailed in Appendex B). This metric provides a comprehensive assessment of UNiCS’s performance across the entire runtime rather than at specific runtime points, and its aggregated nature also helps mitigate the impact of randomness in performance assessment. Then, the optimal $t_{trans}$ with the smallest area is identified. After obtaining training data, a linear model is fitted to these optimal values to derive the transition policy. This simple model is chosen to avoid overfitting while maintaining interpretability and computational efficiency. UNG then uses the trained policy to automatically determine $t_{trans}$ for testing instances.
@@ 4 Computational Studies
Table 1: Relative gap (%) to the BKS on the Large set (10 runs per instance). “-” in Best / Avg columns indicates no solutions produced in 10 runs / solutions produced only in some runs. Bold indicates best performance. “Total” row summarizes average gaps across Large set.
@@
4.1 Experimental Setup
Training UNG. The scoring network SGN within UNG was trained on two-dimensional Euclidean TSP instances with 500 to 1000 nodes, where node coordinates were uniformly sampled from a unit square. Since the number of optimal edges for each instance is proportional to the number of nodes, $\frac{500000}{V}$ instances were generated for each problem size $V$ in the training set, resulting in approximately 346,000 training instances in total. The optimal edges $E^{*}_{o}$ for each instance were obtained using Concorde. Each instance was converted to a sparse directed graph with $\gamma=20$ edges per node (see Section 3.1). The SGN consists of $L=25$ sparse graph convolution layers with a hidden dimension $D=128$ and 4 attention heads. The model was trained for 100 epochs using Adam optimizer with a learning rate of $10^{-4}$.
For the transition policy within UNG, 56 TSP instances ranging from 3,000 to 30,000 nodes were generated using the same uniform distribution described above. For each instance, the transition time was varied from 50 to 650 seconds in 50-second increments. The optimal $t_{trans}$ was then determined by evaluating the area under the optimality gap curve (detailed in Section B in the supplementary material[20]).
TSP Benchmarks. All two-dimensional Euclidean TSP instances with more than 10,000 nodes were collected from three TSP benchmark sets: TSPLib [26], National [8], and VLSI [27], totaling 38 instances. These benchmarks represent diverse real-world scenarios. VLSI instances are derived from integrated circuit applications, National instances are based on city distances in different countries, while TSPLib includes various sources such as logistics, circuit board drilling, and transistor routing. The collected instances were further divided into two sets: Large set (32 instances, 10,000-40,000 nodes) and Ex-Large set (6 instances, 40,000+ nodes). The Large set was further categorized by source: VLSI (21 instances), National (7 instances), and TSPLib (4 instances). The largest TSP instance considered here contains 71,009 nodes.
Competitors and Hyperparameter Settings. UNiCS was compared against state-of-the-art heuristic methods EAX [24] and LKH version 3 [11], as well as hybrid solvers NLKH [32], VSR-LKH [36], and RHGA [37]. To verify the effectiveness of UNG in guiding $AB$-cycle selection, a version of UNiCS (named UNiCS-P) that removes LS phase and uses only PBS phase was also included in the comparison. Essentially, UNiCS-P is a variant of EAX equipped with UNG guidance. For all EAX-based methods (including EAX and UNiCS-P), the EAX-specific hyperparameters were set according to the original EAX paper. All LKH-based methods (including LKH, NLKH, and VSR-LKH) utilized LKH’s default hyperparameter settings. NLKH and UNiCS shared the same SGN architecture and training procedure. For RHGA, hyperparameters were set as specified in the original paper. The probability of random selection of $AB$-cycles in UNiCS, i.e., $\eta$, was set to 0.5 based on the preliminary experiments.
Additionally, a comparison was conducted with LEHD[21] , a state-of-the-art deep learning-based method for TSP. Under same runtime on Medium Set(18 instances, 5,000-10,000 nodes) representative of real-world logistics scenarios, LEHD exhibited a 62.7% gap from BKS, exceeding UNiCS’s performance (0.07%) by over three orders of magnitude. The relative performance gap between LEHD and UNiCS widens by 3.06× when scaling from 5,000 to 10,000 nodes, revealing fundamental limitations in neural architectures for real-world TSP optimization. As these medium-sized instances remain smaller than the instance in Large set, the full detail of this comparison is relegated to Section E in the supplementary material[20].
Testing Methods. To make fair comparison, the testing of all methods was conducted on the same server with two AMD EPYC 7713 CPUs, a NVIDIA A800 80G and 512GB RAM running Ubuntu 20.04, and each method was evaluated using a single CPU core with 10 independent runs per instance. To ensure reproducibility, random seeds started from 42 and were incremented by 60 for each subsequent run. Three runtime budgets were considered in the experiments: short (1800s), medium (3600s), and long (7200s). All reported gaps in the experiments represent the relative gap between the algorithm’s results and the Best-Known Solution (BKS), where BKS refers to the optimal solution computed by exact algorithms or the best solution currently documented in the scientific literature for specific TSP instances. Due to significantly inferior performance compared to other methods, the RHGA results are omitted in the paper and presented in Section A.2 in the supplementary material[20].